### v1

# Gen-C: Populating Virtual Worlds with Generative Crowds

Abstract
Over the past two decades, researchers have made significant advancements in simulating human crowds, yet these efforts largely focus on low-level tasks like collision avoidance and a narrow range of behaviors such as path following and flocking. However, creating compelling crowd scenes demands more than just functional movement—it requires capturing high-level interactions between agents, their environment, and each other over time. To address this issue, we introduce Gen-C, a generative model to automate the task of authoring high-level crowd behaviors. Gen-C bypasses the labor-intensive and challenging task of collecting and annotating real crowd video data by leveraging a large language model (LLM) to generate a limited set of crowd scenarios, which are subsequently expanded and generalized through simulations to construct time-expanded graphs that model the actions and interactions of virtual agents. Our method employs two Variational Graph Auto-Encoders guided by a condition prior network: one dedicated to learning a latent space for graph structures (agent interactions) and the other for node features (agent actions and navigation). This setup enables the flexible generation of dynamic crowd interactions. The trained model can be conditioned on natural language, empowering users to synthesize novel crowd behaviors from text descriptions. We demonstrate the effectiveness of our approach in two scenarios, a University Campus and a Train Station, showcasing its potential for populating diverse virtual environments with agents exhibiting varied and dynamic behaviors that reflect complex interactions and high-level decision-making patterns.
Keywords crowd simulation, data-driven method, multi-agent navigation, crowd authoring, variational graph auto-encoder
Figure 1: Framework Overview: We leverage a Large Language Model to generate high-level crowd scenarios, generalize them into synthetic crowd data encoded as graphs, and learn a graph-feature space. Using input textual conditions, we sample this space to generate novel crowd scenarios.
1 Introduction
To facilitate digital human creation, the field of computer graphics has experienced a dramatic increase in the number of tools, approaches and algorithms focusing on creating compelling crowd motions [1, 2]. Currently, several game engines and companies specializing in offline modeling and rendering offer tools to author crowd simulations [3, 4, 5, 6]. Nevertheless, a lot of manual work is still needed by the animators and level designers along with deep familiarity with the underlying tools in the production workflow in order to obtain desired simulation results. While a number of automated solutions to crowd simulation have been proposed over the past decade, they mostly focus on low-level tasks like resolving collisions between agents and a narrow range of behaviors such as steering and rouping [7]. In contrast, real humans are capable of exhibiting high-level behaviors of enormous complexity and diversity based on their own goals and desires and the affordances from the environment. For example, we stop on the street to briefly talk to friends, browse through shopping windows, buy a ticket before heading to the train platform, etc. Currently, such behaviors are either absent from existing simulators resulting in "boring" and repetitive simulations or are scripted which makes them difficult to scale.
In this paper, we introduce Gen-C, a generative framework that allows users to automatically populate virtual worlds with diverse crowds. At the core of our approach lies a crowd scenario graph, which is a time-expanded graph that encodes the spatial and temporal evolution of the crowd in a scene, capturing high-level agent-to-agent and agent-to-environment interactions over time. With the advent of big data and deep learning, an obvious solution would be to learn such crowd scenario graphs from real crowd data. Indeed, a number of data-driven approaches have been proposed that learn crowd models from trajectories extracted from real-world crowd data [8, 9, 10, 11]. Nevertheless, the dependency on having to collect and analyze real crowd data can quickly become a barrier, limiting also the generalization capabilities of such solutions.
To break free from the dependency on real crowd data, we propose to automatically synthesize interaction datasets by leveraging the power of large language models (LLMs) such as the GPT framework [12]. As querying LLM APIs can become prohibitively expensive, we use them to create a few indicative crowd scenarios that describe agent-agent and agent-environment interactions for given input scenes, which we subsequently expand and generalize through simulations to construct corresponding crowd scenario graphs. Given such graphs, a generative model is trained to learn a latent space of agent behaviors and interactions and automate the synthesis of new graphs. As most existing graph embedding algorithms are mainly developed for static graphs, they cannot capture the complex and dynamic nature of a crowd scenario graph. To that end, we propose a novel model architecture based on variational graph autoencoders (VGAEs) [13] that can learn how agents interact with each other and how they behave with respect to their environment. The trained model can be conditioned on natural language, allowing users to automatically author scenes with diversified crowds from text input.
Our approach is inspired by recent trends in generative AI. We note that a lot of work in computer graphics and vision has applied generative models for text-driven human motion synthesis [14] with notable success. Despite being capable of synthesizing diverse behaviors, such works typically focus on individual humanoids.
More closely to our domain, generative models have also been explored to steer groups of agents [15] as well as to synthesize crowd trajectories [16] by leveraging guided diffusion. Our work is complementary to such approaches, as we seek to learn the high-level behaviors of crowds that can facilitate the generation of such trajectories. By leveraging a generative model, we strive for an approach that can generalize to different settings and number of agents, allowing characters to exhibit an array of high-level behaviors.
Overall, we propose:
  1. 1.
    Gen-C, a framework for text-guided authoring of high-level crowd behaviors that
  2. 2.
    breaks free from the dependency on real-world crowd data via the use of LLMs and simulations through
  3. 3.
    a text-conditioned graph generative model that relies on a dual VGAE architecture to encode the spatial and temporal evolution of agent behaviors, interactions, and movement.
2 Related Work
2.1 Authoring Virtual Crowds
Controlling crowd simulations at a higher level is crucial to several scenarios since it allows users to easily and efficiently author agents’ behaviors according to their wishes. This requires intuitive tools which are highly dependent on what simulation aspect the users aim to control [2]. Early approaches to authoring large crowds included the use of pre-computed collision free trajectories in the form of crowd patches [17, 18] or crowd animations that could be connected together to form larger simulations  [19]. Others, devised user interactive tools to allow for the definition of navigation fields in pre-defined environments [20]. For authoring navigation behavior in crowd scenes, past literature heavily uses sketch-based interfaces [21, 22]. In contrast, authoring the animation and visualization aspects usually entails asset and template manipulation [23, 24]. Editing has also been widely explored in literature, with the more user-friendly systems incorporating manipulation handles. For example, Kwon et al. [25], propose deformation gestures as a post-processing tool for group motion editing. One of the most challenging aspects in authoring is controlling high-level behaviors such as describing agendas, desires and even personalities; these behaviors benefit the expressiveness and realism of simulations [26]. Complimentary to these approaches, we propose a text-driven generative approach to automatically author high-level crowd behaviors in specific environments, capitalizing on the implicit knowledge of LLMs for human decision making in different urban spaces.
2.2 Heterogeneity in Virtual Crowds
Generating behavioral diversity in virtual crowds has been addressed in the literature with a variety of approaches; early works use predefined rules for different behaviors to achieve such heterogeneity [27, 28]. Ren et al.  [29] incorporated diverse properties in groups of agents controllable via user constraints, while more generically several researchers [30, 31, 32] focused on behavior diversity in terms of personality traits; these expose some parameters to users to allow for control of agent behaviors. Data-driven approaches have also been prominent in the literature, where the crowd simulation model is implicitly defined by example data. Interestingly, early data-driven approaches were graph-based [33, 25], similar to early methods in the character animation literature [34]. These methods however remain limited to group navigation (e.g., flocking) and fail to reflect the variations of behaviors found in human crowds. The need to simulate behavior diversity found in real-life crowds, led to more sophisticated and practical approaches, including approaches that learn crowd models from trajectories extracted from real-world crowd data [35, 36, 37], including works that define secondary high-level actions for increased realism [38, 39], blending different crowd styles represented by the input data [21], and more recent solutions like deep reinforcement learning [40], and even in combination with imitation learning [16, 11]. Despite the significant improvements achieved by these methods, the results highly depend on the variability and amount of input data. Contrary to all of these works, Gen-C defines heterogeneity through the use of crowd scenario graphs that are automatically generated through LLM queries and a basic simulator allowing us to move away from the dependency on collecting real crowd data.
2.3 Large Language Models and Virtual Agents
Large language models have shown to be effective to domains beyond natural language processing such as task planning, embodied reasoning, autonomous agents, and floorplan design [41, 42, 43, 44]. Here, we explore an architecture that allows a LLM to generate a sequence of crowd behaviors over time and space that can then be generalized through simulations to create synthetic crowd data. We further exploit the power of generative NLP models to enable text-guided synthesis of high-level crowd behaviors. Along these lines a lot of work has explored generative models for text-driven human motion generation tasks [14]. Despite impressive results, including generating scene-and-language conditioned motions [45, 46, 47, 48, 49], such solutions typically focus on generating animations for an individual character. But a human crowd is much more than that, consisting of multiple agents performing a sequence of behaviors while interacting with each other and the environment. Along these lines, the recent work of [15] uses a text-based diffusion model to steer crowds of agents but does not extend to other types of behaviors. Our work seek to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
2.4 Graph Generation
Besides our domain, learning graph representations is a fundamental task for many real-world applications [50]. While a lot of prior work has focused on learning static graphs with deep learning models such as VGAEs [13], recent techniques have explored learning of dynamic graphs where the graph structure itself changes as time progresses, including neural temporal point processes and attention mechanism for modeling time-dependent events [51], and deep architectures such as Temporal Graph Networks [52] and Spatio-Temporal Graph Convolutional Networks [53] that can capture temporal dependencies and predict future states of evolving graphs. While highly relevant to our work, we focus on crowd scenario graphs: highly dynamic graphs with complex features that capture the evolution of human decision-making. These graphs are more intricate than the time-evolving graphs studied in prior work, which are typically designed for tasks like link prediction, node classification, recommendation systems, and social network analysis. Closely related, conditional generation of graphs has also gained significant attention, with majority of works focusing on generating the graph structure only while the node and edge features are predetermined as compared to our approach. For example, NGG [54] uses conditioned latent diffusion, GraphVAE [55] integrates conditional codes for guided generation, and MOOD [56] guides a diffusion sampling process with property gradients. Sequential models like in [57] utilize conditions to initialize node and edge creation, while others incorporate conditional codes into message-passing schemes [58]. Complementary to these works , our Gen-C framework embeds both graph structures and node features into two distinct, prior-conditioned latent distributions, enabling their generation based solely on conditional text input.
3 Gen-C Overview
Preliminaries. We define a High-Level Scenario as the description of actions, interactions, and environment structure. Action is defined as a label from $\mathcal{A}ct=~{}${stand still, sit, wait, wander, queue, object interact, talk, meet, enter/exit, leave group, talk to phone, wave at, read, look at, carry}, that guides the agent’s behavior. When multiple agents perform a shared action, e.g. talk, we refer to that as an Interaction. Environment Structure defines the position, scale, and orientation of various locations; locations can be abstract, however we categorize them in: $\mathcal{L}oc=~{}${building, room, entrance, exhibit, furniture, outdoor area, item, service area}. To simplify our problem, we assume a predefined list of actions and location categories, and an action is always paired with a location, e.g. {talk, room}.
Gen-C Framework. Our proposed framework for populating virtual environments with generative crowds consists of three main components (Fig. 1): Synthetic Data Generation, Learning, and Crowd Scenarios Generation. To generate the synthetic data, we first use a LLM to create diverse crowd scenarios through targeted queries that capture context, environment placement, action probabilities, and movement probabilities. These scenarios are extended using a preliminary simulator to introduce randomness and generalize the data, which is then transformed into graph-like structures representing action sequences, relationships, and interactions. Given such graphs, we train two VGAE models: one learns to reconstruct the graph structures capturing agent interactions, while the other focuses on node features representing agent actions and navigation. A text-conditional network acts as a prior, integrating textual information to ensure the generated graphs align with given descriptions. The trained models can then synthesize novel crowd scenarios by conditionally sampling from the learned latent distributions, generating contextually aligned scenarios.
4 Synthetic Data Generation
Simulating diverse and interesting high-level crowd behaviors is a complex task that relies heavily on high-quality data. However, collecting and annotating real-world crowd data is a highly challenging and resource-intensive task. We propose to overcome this limitation by leveraging recent advancements in LLMs to generate synthetic crowd data that adhere to realistic rules and behaviors.
4.1 Querying Crowd Scenarios
To generate simulations that faithfully reflect crowd dynamics, we use a series of tailored LLM queries, each designed for a specific task. The process begins with a single sentence input $S_{in}$ describing a crowd scenario, which acts as the initial seed for the generation of a scenario (see top-left in Fig. 1 for an example input). Specifically, we select OpenAI’s gpt-4o [59] model and run four sequential queries as presented below:
Expanding the Scenario ($Q1$).
The first query requests the LLM to expand $S_{in}$ into a detailed paragraph $Q1_{par}$ that describes the scenario. This enriched description serves as the foundation for further steps as it includes individual agent behaviors and interactions, while also provides information about the environment context.
Defining the Environment ($Q2$).
Using the detailed description $Q1_{par}$ and a randomly selected environment size, the second query generates a meaningful environment layout $Q2_{env}$ specifying locations and areas relevant to the scenario. For example, the output might define places such as a “coffee shop”, “park”, or “entrance area”, along with their position, scale and orientation.
Assigning Location-Specific Actions ($Q3$).
For each generated location in $Q2_{env}$, the third query identifies a list of plausible actions from $\mathcal{A}ct$ that agents might perform in that area. For instance, in a “coffee shop,” actions like “queue,” “meet,” or “sit” are expected. The LLM takes as input the $Q1_{par}$ and $Q2_{env}$ and is requested to select the top-5 actions that are most suitable for each area and assign selection probabilities to them; we ensure the total probability sums to 1. We note that we allow multiple possible actions per location to promote diversity in the generated simulations.
Building Movement Plans ($Q4$).
The fourth query focuses on defining how agents move between different locations within the environment. This step is similar to $Q3$, however, assigns movement probabilities from one location to every other in the environment, creating a dynamic movement plan for the agents.
Finally, we combine the responses from all four queries and form a high-level crowd scenario $S$, generated by a high-level textual description $S_{in}$.
4.2 Collecting Individual Behaviors
After gathering multiple crowd scenarios from the LLM, we need a mechanism to utilize them so we can generate individual actions and interactions for each agent. Thus, we construct a simple simulator in Unity, where we parse each scenario $S$ and run multiple simulations.
For each LLM scenario, we build the environment as described by $Q2_{env}$. Then, we randomly spawn agents in the environment, either as individuals or in small groups of two or three (all agents in a group perform the same action). Initially, each agent or group, queries its current location action list and randomly selects a future action based on the assigned probabilities; each action has its own duration based on the context. After completing the action, the agent or group of agents selects the next location based on the movement probabilities of its current location. This process repeats in cycles, with data collected over a period of 2 to 3 minutes.
To introduce variability and prevent agents from consistently selecting actions and movements with the highest probabilities, we use a temperature parameter $\text{Temp}\in[.7,1]$ to amplify lower probabilities. Temp is randomly initialized and gradually increased to 1 by the end of the simulation.
Finally, for each agent, we record its action sequence throughout the simulation. Each entry in the sequence captures the selected action and its location. In case the current action is shared with others (interaction), we include the ID of other agent/s in the entry. This detailed record $Rec_{\alpha_{i}}$ for each agent ${\alpha_{i}}$, provides a comprehensive view of the agent’s behavior and interactions within the simulated environment over its entire lifespan.
4.3 Crowd Scenario Graph Representation
In order to effectively train a generative crowd model, it is essential to select an appropriate data structure that can accurately represent the sequence of actions and interactions of agents. The chosen structure must be capable of: (a) capturing interactions between individual agents, and (b) representing their sequence of actions over time. We propose using a graph structure as it fulfills these requirements. In particular, we introduce the concept of crowd scenario graph, which is a a time-expanded graph that encodes both the actions and interactions of agents over a fixed temporal horizon, modeling the dynamic behaviors and relationships inherent in crowd scenarios.
Figure 2: Structure of a crowd scenario graph. Each node contains an action label, a location category, an agent ID, and timestep as $\alpha_{i}$_$t$.
As described in Section 4.2, during simulation, for each agent ${\alpha_{i}}$, we collect $Rec_{\alpha_{i}}$. We iterate through $Rec_{\alpha_{i}}$ and create a node $V_{\alpha_{i}}^{t}=({\alpha_{i}},A_{\alpha_{i}}^{t},L_{\alpha_{i}}^{t})$, where $t$ is the current timestep, $A_{\alpha_{i}}^{t}\in\mathrm{A}ct$ and $L_{\alpha_{i}}^{t}\in\mathrm{L}oc$. We define the sequence of actions of ${\alpha_{i}}$ using a “sequence” edge $E_{\alpha_{i}}^{t}=(V_{\alpha_{i}}^{t},V_{\alpha_{i}}^{t-1})$. Next, we process the interactions between agents by connecting nodes from different agents that share an action at a given timestep. We create a “share edge” $E_{{\alpha_{i}},{\alpha_{j}}}^{t}=(V_{\alpha_{i}}^{t},V_{\alpha_{j}}^{t})$ for each agent ${\alpha_{i}}$ interacting with ${\alpha_{j}}$ at timestep $t$. Fig. 2 presents a simple example of a crowd scenario graph. At this stage, various sequences of nodes may become connected, forming subgraphs that represent the interactions of agents within a group. In cases where an agent has no interactions with others, the initial set of the agent’s nodes still forms a subgraph on its own.
Thus, for each simulated scenario $S$, we construct an undirected graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the edge set, comprising a collection of subgraphs $SG$. Finally, we treat each entry of $SG$ as an individual sample $(S_{in},SG_{i},agents_{i})$, where $S_{in}$ is the textual-description of the current scenario, $SG_{i}$ is the subgraph of group $i$, and $agents_{i}$ is the number of agents in current group.
5 Learning crowd scenario graphs
While a synthetic crowd scenario graph can be used to drive a crowd of simulated agents, ideally we want to be able to generalize without having to manually compute a new graph from scratch every time the input conditions change. To do so, given a dataset consisting of crowd scenario graphs, we will employ a generative model that can reason about the graphs and generate new ones.
5.1 Data Preprocessing
Before training, we apply some preprocessing steps to prepare our data. For each sample $(S_{in},SG_{i},agents_{i})$, we treat subgraph $SG_{i}$ as an individual undirected graph $G_{i}=(V_{i},E_{i})$, where $n=|V_{i}|$ (number of nodes) and $m=|E_{i}|$ (number of edges). The adjacency matrix $\mathcal{A}_{G_{i}}\in\mathbb{R}^{n\times n}$ of a graph $G_{i}$ is a symmetric matrix that encodes edge information in the graph. The value of the $i^{th}$ row and $j^{th}$ column is equal to 00 if there is no edge between $v_{i}$ and $v_{j}$. Otherwise, it set to $1$ or $-1$, if edge’s type is “sequence” or “share” respectively.
Node Ordering. First, we design a specific node ordering scheme to ensure a consistent graph representation. The adjacency matrix of a graph can be constructed with different permutations of node indices, which can lead to inconsistency and increased complexity during model training, especially for graph reconstruction tasks. To address this, we define a canonical ordering for the nodes of each graph based on the agent ID (${\alpha_{i}}$) and $t$. We sort all nodes of a graph using ${\alpha_{i}}$ and $t$ as a primary and secondary key respectively. For instance, $\{V_{0}^{1},V_{1}^{0},V_{1}^{1},V_{0}^{0}\}$ will be transformed to $\{V_{0}^{0},V_{1}^{0},V_{0}^{1},V_{1}^{1}\}$. Fig. 3 presents a comparison between the adjacency matrices using an abstract and a canonical ordering for different numbers of nodes.
Figure 3: Canonical Ordering (top) vs Abstract Ordering (bottom). In canonical, nodes are sorted first by agent ID, then by the node’s sequence in agent’s action, ensuring consistent graph representation.
Node Features. We annotate the graphs nodes using feature vectors. The feature vector of each node includes: the action, location, unique agent id, sequence index and interaction status. First, we use action and location as two one-hot vectors (sizes 15 and 8) from the available actions $\mathcal{A}ct$ and available location categories $\mathcal{L}oc$. For agent ID (${\alpha_{i}}$) and timestep $t$ in sequence, we set the maximum number of agents per subgraph equal to 6 and the maximum sequence length to 10; these choices reduce the complexity of generated graphs.
We construct another two one-hot encodings (sizes 6 and 10) for them too. We encode the interaction status of the current node, indicating whether the current action is shared or not with another one-hot vector (size 2). To generate the final node features, we use embedding layers to ensure that node features are dense and learnable representations suitable for downstream tasks. Each encoding is then passed through its respective embedding layer: actions ($\mathbb{R}^{15\times 6}$), locations ($\mathbb{R}^{8\times 4}$), agents ($\mathbb{R}^{6\times 5}$), timestep ($\mathbb{R}^{10\times 4}$), and interaction status ($\mathbb{R}^{2\times 2}$). The embeddings are concatenated to form the final node feature representation $\mathbf{X}\in\mathbb{R}^{n\times d}$ where $d$ is the feature dimension size. Additionally, we include the top 4 eigenvectors of the normalized graph Laplacian as node features to encode global structural properties [60]. Thus, the final node feature dimension becomes $d=21+4=25$.
Textual Condition. We convert $S_{in}$ to textual condition by first detecting and emphasizing verbs in the text to highlight action-relevant information; we utilize the spaCy library [61]. Each $S_{in}$ is augmented by prefixing it with a list of extracted verbs, and its embedding is generated using the pretrained Sentence Transformer model all-MiniLM-L6-v2 [62].
Figure 4: Training Model Architecture. We combine two VGAEs, one for graph reconstruction and the other for node feature reconstruction. Each model’s encoder compresses high-dimensional inputs into latent representations, while a shared Condition Network serves as a prior to regularize the latent spaces.
5.2 Training
We train two separate conditional VGAE models and learn two distinct latent spaces, one for reconstructing graph structure (VGAE-S) and the other for reconstructing node features (VGAE-F). Each model consists of an Encoder, a Decoder, and a Condition Net. Figure 4 illustrates the training model architecture of our framework.
Graph Structure and Features Encoder. Both VGAE-S and VGAE-F models use the same encoder structure to map input graphs into latent representations $Z_{S}$ and $Z_{F}$, respectively, by taking as input the adjacency matrix $\mathcal{A}_{G}$ and features $\mathbf{X}_{G}$. We employ a sequence of GINE layers [63], a variant of message-passing neural networks enabling the use of edge features, to iteratively update node embeddings based on their neighbors and edge features. We note that before the encoding, self-loops are added to ensure their own features contribute to the updated representation during message-passing. Each layer aggregates node features and edge features as:
$\mathbf{h}_{v}^{(k)}=\text{MLP}^{(k)}\big{(}\mathbf{h}_{v}^{(k-1)}+\textstyle\sum_{u\in\mathcal{N}(v)}e_{uv}\cdot\mathbf{h}_{u}^{(k-1)}\big{)},$ (1)
where $\mathbf{h}_{v}^{(k)}$ represents the node feature at layer $k$, $\mathcal{N}(v)$ denotes the neighbors of node $v$, and $e_{uv}$ are the edge features. Each MLP is composed of linear layers, LeakyReLU activations, and BatchNorm layers. We list specific choices for the networks in our experiments in Section 7. After the message-passing steps, the node embeddings are pooled into a single graph-level embedding using global mean pooling. The pooled embedding is then concatenated with the conditional vector $\mathbf{C}$ and is passed through a fully connected layer with LeakyReLU activation to produce the mean $\mu$ and standard deviation $\sigma$ of a Gaussian distribution. This enables the encoder to parameterize the posterior distribution $q(Z|X,C)$.
Graph Structure Decoder. Reconstructs the padded graph adjacency matrix $\mathcal{A}_{G^{\prime}}$ from $Z_{S}$; we set the maximum number of nodes equal to 40. It uses a MLP to transform the concatenated latent and conditional representations into a flattened adjacency matrix. An MLP maps the concatenated latent and conditional representations to the upper triangular part of the adjacency matrix, which is symmetrized to produce the full matrix.
Features Decoder. Reconstructs node features $\mathbf{X}_{G^{\prime}}$ from $Z_{F}$, and $\mathcal{A}_{G}$. The decoder employs a series of GINE layers, where the input to each layer is a combination of the expanded $Z_{F}$, $\mathbf{C}$. Each GINEConv layer applies a message-passing operation, followed by non-linear transformations using LeakyReLU activations and LayerNorm for stability. The final layer projects the hidden representations into the node feature space, with size 23.
Condition Net. Both VGAE models use the same condition network structure that encodes the embeddings of text ($S_{in}$) and number of agents to produce the condition vector $C$. We utilize $\mathbf{C}$ to train two Prior Networks that parameterize the prior distributions $p(Z|C)$ of the latent variables of VGAE-S and VGAE-F, computing their respective means $\mu_{\text{prior}}$ and standard deviations $\sigma_{\text{prior}}$. Additionally, $C$ is integrated into both the encoding and decoding phases. In the encoder, it is concatenated with the graph-level features embedding to produce the posterior distribution $q(Z|X,C)$, parameterized by $\mu$ and $\sigma$., while during decoding it guides the reconstruction by conditioning the latent space.
During training, for each textual sentence $S_{in}$ we prepare a list of 20 paraphrased versions and randomly select one at each epoch to enhance generalization. For both VGAEs we seek to maximimze the evidence lower bound that results in a reconstruction loss and a KL divergence loss. The graph structure reconstruction loss $\mathcal{L}_{S}$ is calculated by comparing $\mathcal{A}_{G^{\prime}}$ with the input matrix $\mathcal{A}_{G}$; note that the edge type (“sequence” or “share”) is predicted too. The node feature reconstruction loss $\mathcal{L}_{F}$ is computed by comparing the action and location category class for each node between $\mathbf{X}_{G^{\prime}}$ and $\mathbf{X}_{G}$. Formally, $\mathcal{L}_{S}$ and $\mathcal{L}_{F}$ are calculated as:
$\mathcal{L}_{S}$ $=\sum_{i,j}\text{SmoothL1}(\mathcal{A}_{G_{ij}},\mathcal{A}_{{G^{\prime}}_{ij}}),and$ (2)
$\mathcal{L}_{F}$ $=\sum_{i=1}^{N}\left(\text{CELoss}(\mathbf{X}_{G_{i}}^{\text{act}},\mathbf{X}_{{G^{\prime}}_{i}}^{\text{act}})+\text{CELoss}(\mathbf{X}_{G_{i}}^{\text{loc}},\mathbf{X}_{{G^{\prime}}_{i}}^{\text{loc}})\right),$
where SmoothL1 denotes a smooth L1-error and CELoss denotes the cross-entropy loss. The KL divergence loss $\mathcal{L}_{KL}$ measures the difference between the posterior distribution $q(Z|X,C)$ and the prior distribution $p(Z|C)$, and is computed by:
$\mathcal{L}_{\text{KL}}=\mathbb{E}_{z\sim q(Z)}\left[\log q(Z\mid X,C)-\log p(Z\mid C)\right].$ (3)
[…]
6 Text-Conditioned Scenario Generation
During inference, the learned latent spaces of our model enable the generation of novel crowd scenarios guided by a textual condition. Specifically, graph structures can be sampled from the prior distribution $p(Z_{PS}|C)$, while their node features are populated using the node feature prior $p(Z_{PF}|C)$. This approach ensures that both the structure and features of the generated graphs are contextually aligned with the given a condition $C$. Each generated graph represents a subgraph that captures the interactions and actions of a single group of agents, as detailed in Section 4.3. As the number of agents per group is defined in $C$, this makes our framework inherently scalable, enabling the generation of an arbitrary number of agents or groups of agents for any given input condition.
We parse the generated graphs to collect a sequence of actions-locations and interactions per agent. Then, to generate a simulation, a simulation environment can be either (a) manually provided by the user or (b) generated with a simple LLM query similar to $Q2$, as described in Section 4.1. Given the environment, we run simulations of the generated scenarios in Unity [5]. We provide visual results in Fig. 7, while related animations are also presented in the supplementary video.
7 Experiments and Evaluation
We train our model on two synthetic datasets representing crowd scenarios for a University Campus and a Train Station. For each dataset, we begin by defining 100 unique input sentences describing crowd scenarios; an example input is “Outside the station, passengers sit on benches to read a newspaper and talk on the phone.” Using the process detailed in Section 4.1, we generate 5 variations for each sentence, resulting in 500 crowd scenarios per dataset. Graph datasets are then constructed as outlined in Section 4.2 and Section 4.3, with the University Campus dataset containing $135k$ subgraphs and the Train Station dataset $110k$. The training data and models will be released upon acceptance. We split each dataset into 75% for training, 15% for validation, and 10% for testing, and train a separate model per dataset. For model training (cf. Fig. 4), we use a graph encoder consisted of 2 GINE layers with a hidden dimension of 96, each followed by a dropout layer with a rate of .2. Both VGAE-S and VGAE-F have a latent dimension of 16. The graph structure decoder uses 3 linear layers with a hidden dimension of 128, applying a tanh activation after the final layer. The feature decoder employs 3 GINE layers with a hidden dimension of 128, followed by a layer mapping to a 23-dimensional feature space. The condition network projects text embeddings from 384 to 128 dimensions and agent embeddings from 32 to 128, combining them into a conditional vector of size 128. The prior network maps this conditional vector to 16 dimensions using 2 linear layers. LeakyReLU activations are used throughout, with BatchNorm in the encoder and LayerNorm in both decoders for stability. The models are trained for 2,000 epochs with a batch size of 256, a learning rate of 0.001, and the Adam optimizer with a weight decay of 0.0004. Every cycle for the $\beta$ annealing lasts 200 epochs. Training each model required approximately 4 hours on a single NVIDIA RTX 4070 Ti GPU.
Quantitative Results.
We test the reconstruction quality of our model relying solely on the conditional prior $p(Z|C)$. Specifically, we employ the test set for each dataset and sample scenarios (crowd scenario graphs) using the $S_{in}$ of each ground-truth sample. We employ various metrics to evaluate reconstruction quality. We aim to evaluate both the reconstruction of (a) graph structure and (b) features. For (a) we use a list of well-known graph metrics [65] and focus both on local and global structure. These metrics are: i) Node degree, which measures the number of connections a node has, capturing the graph’s local connectivity; ii) Clustering coefficient that quantifies the tendency of nodes to form tightly connected groups; iii) Graph diameter that represents the longest shortest path between any two nodes, reflecting the graph’s overall size; iv) Average path length that calculates the mean shortest distance between all node pairs, assessing the graph’s overall connectivity. For (b), we calculate the frequency of each action and location label from $\mathcal{A}ct$ and $\mathcal{L}oc$.
We refer to  Fig. 5 for corresponding results. For each statistic we calculate the KL Divergence (KLD) between the true and predicted distribution; lower KLD value indicates better alignment between the two distributions. The results demonstrate that our model can generate graphs with similar structural characteristics to the ground-truth data, indicating good alignment between prior and posterior latent space of VGAE-G model. Similarly, the VGAE-F model demonstrates efficiency in feature reconstruction, as the generated labels closely match the data distribution. Our models also perform consistently across datasets with differing dynamics. For instance, in the Train Station dataset, the "carry" label is more frequent, while "wander" is less common compared to the University Campus dataset, reflecting travelers often carrying bags and students wandering in outdoor areas.
We also evaluate the diversity of feature generation in our model. Using the same textual description $S_{in}$, we sample 1k variables from the latent space of VGAE-F and collect each label’s frequency. The results for four user-input scenarios are presented in Fig. 6, demonstrating that the model produces diverse outputs rather than being limited to a fixed set. Moreover, the generated results remain consistent with the input condition $S_{in}$.
Ablation Study.
In Table 1 we present the KL Divergence between the true and predicted samples using the graph structure and feature metrics presented above for two variations of our model: (a) without using the custom canonical node ordering (w/o Cano.), detailed in Section 5.1, and (b) using a single latent space for both graph structure and features (Single). The results demonstrate the effectiveness of the proposed canonical order scheme, as its absence leads to poorer performance in graph reconstruction. Additionally, using a single latent space significantly affects feature reconstruction (action and location), highlighting the advantage of employing separate latent spaces to better capture the data’s complexity.
Qualitative Results.
In Fig. 7 we provide rendered results generated from different textual conditions for both datasets. Note that the selected $S_{in}$ can be abstract, however we keep them similar to the context of each dataset. We encourage readers to refer to the provided supplementary video for more detailed results.
Table 1: Ablation Study. Comparing our Gen-C training routine to training without canonical node ordering (w/o Cano.) and with using a combined VGAE model for both graph structure and node features (Single). Reported numbers denote the KL divergence (KLD) between ground truth and predicted samples on our testing datasets (lower is better).
Metrics University Campus Train Station
w/o Cano. Single Gen-C w/o Cano. Single Gen-C
Degree .2578 .0053 .0045 .4710 .0045 .0055
Clustering Coeff. .5143 .0203 .0046 .4511 .0096 .0074
Diameter .1996 .1425 .1200 .2642 .0369 .0820
Avg. Path Length .1877 .2244 .0360 .6330 .1005 .0500
Action .4149 .3009 .0640 .2699 .3078 .1100
Location .0925 .0316 .0027 .0270 .0480 .0088
8 Discussion and Future Work
We propose Gen-C, a framework designed to populate virtual environments with multiple agents exhibiting diverse high-level actions and interactions, conditioned on textual input. Our approach leverages an LLM and two learned latent spaces to enable the generation of crowd scenario graphs that drive agent behaviors.
Overall, our system generate plausible simulations, allowing the automatic creation of virtual environments with agents exhibiting diverse, high-level actions beyond typical low-level crowd behaviors like goal-seeking and collision avoidance. However, several limitations exist. The simulator used to expand LLM-generated scenarios relies on probabilistic action sequences, thus the sequence of actions may not be optimal. Agents cannot switch actions mid-execution, and action durations are manually defined. Additionally, behaviors are constrained to a predefined list of high-level actions, requiring retraining to accommodate new actions.
For future work, we aim to integrate our framework with a crowd simulation framework, allowing high-level behaviors to control low-level navigation policies, creating more dynamic and realistic simulations. Additionally, we plan to explore combining latent spaces from different datasets, such as "indoor" and "outdoor" scenarios, using techniques such as distillation to enable lifelong learning and dynamic simulation adaptability. We envision that these will allow us to populate virtual worlds with diverse human-like characters, paving the road towards automatic generation of digital crowds that can be applied to a variety of important areas.

### v2

@@
 Abstract
-Over the past two decades, researchers have made significant advancements in simulating human crowds, yet these efforts largely focus on low-level tasks like collision avoidance and a narrow range of behaviors such as path following and flocking.
-However, creating compelling crowd scenes demands more than just functional movement—it requires capturing high-level interactions between agents, their environment, and each other over time.
-To address this issue, we introduce Gen-C, a generative model to automate the task of authoring high-level crowd behaviors.
-Gen-C bypasses the labor-intensive and challenging task of collecting and annotating real crowd video data by leveraging a large language model (LLM) to generate a limited set of crowd scenarios, which are subsequently expanded and generalized through simulations to construct time-expanded graphs that model the actions and interactions of virtual agents.
-Our method employs two Variational Graph Auto-Encoders guided by a condition prior network: one dedicated to learning a latent space for graph structures (agent interactions) and the other for node features (agent actions and navigation).
-This setup enables the flexible generation of dynamic crowd interactions.
-The trained model can be conditioned on natural language, empowering users to synthesize novel crowd behaviors from text descriptions.
-We demonstrate the effectiveness of our approach in two scenarios, a University Campus and a Train Station, showcasing its potential for populating diverse virtual environments with agents exhibiting varied and dynamic behaviors that reflect complex interactions and high-level decision-making patterns.
+Over the past two decades researchers have made significant advancements in simulating human crowds, yet these efforts largely focus on low-level tasks like collision avoidance and basic behaviors such as path following and flocking.
+Creating realistic crowd scenes, however, requires modeling richer, high-level interactions among agents, their environment, and each other over time.
+We introduce Gen-C, a graph-based generative model that shapes high-level agent actions and interactions.
+To bypass the labor-intensive and challenging task of collecting and annotating real crowd video data, we leverage a Large Language Model (LLM) to generate a dataset of crowd scenarios.
+Each scenario is then represented as a set of time-expanded graphs modeling agent actions and interactions.
+Gen-C employs a dual Variational Graph Autoencoder (VGAE) architecture to jointly learn graph structure (agent interactions) and features (agent actions), enabling the population of virtual scenes conditioned on natural language.
+We demonstrate the effectiveness of our approach across diverse scenarios, including a University Campus and a Train Station, showcasing its ability to populate virtual environments with agents that exhibit dynamic behaviors, complex interactions, and high-level decision-making.
 Keywords crowd simulation, data-driven method, multi-agent navigation, crowd authoring, variational graph auto-encoder
@@
 Nevertheless, a lot of manual work is still needed by the animators and level designers along with deep familiarity with the underlying tools in the production workflow in order to obtain desired simulation results.
-While a number of automated solutions to crowd simulation have been proposed over the past decade, they mostly focus on low-level tasks like resolving collisions between agents and a narrow range of behaviors such as steering and rouping [7].
+While a number of automated solutions to crowd simulation have been proposed over the past decade, they mostly focus on low-level tasks like resolving collisions between agents and a narrow range of behaviors such as steering and grouping [7].
 In contrast, real humans are capable of exhibiting high-level behaviors of enormous complexity and diversity based on their own goals and desires and the affordances from the environment.
 For example, we stop on the street to briefly talk to friends, browse through shopping windows, buy a ticket before heading to the train platform, etc.
-Currently, such behaviors are either absent from existing simulators resulting in "boring" and repetitive simulations or are scripted which makes them difficult to scale.
+Currently, such behaviors are either absent from existing simulators resulting in "boring" and repetitive simulations or are manually scripted which makes them difficult to scale.
 In this paper, we introduce Gen-C, a generative framework that allows users to automatically populate virtual worlds with diverse crowds.
@@
 To break free from the dependency on real crowd data, we propose to automatically synthesize interaction datasets by leveraging the power of Large Language Models (LLMs) such as the GPT framework [12].
-As querying LLM APIs can become prohibitively expensive, we use them to create a few indicative crowd scenarios that describe agent-agent and agent-environment interactions for given input scenes, which we subsequently expand and generalize through simulations to construct corresponding crowd scenario graphs.
-Given such graphs, a generative model is trained to learn a latent space of agent behaviors and interactions and automate the synthesis of new graphs.
+As querying LLM APIs can become prohibitively expensive, we use them to create a few indicative synthetic crowd scenarios that describe agent-agent and agent-environment interactions for given input scenes, which we subsequently represent as crowd scenario graphs.
+Given such graphs, a generative model is trained to learn a latent space of agent behaviors and interactions, that automates the synthesis of new graphs (crowd scenarios).
 As most existing graph embedding algorithms are mainly developed for static graphs, they cannot capture the complex and dynamic nature of a crowd scenario graph.
-To that end, we propose a novel model architecture based on variational graph autoencoders (VGAEs) [13] that can learn how agents interact with each other and how they behave with respect to their environment.
-The trained model can be conditioned on natural language, allowing users to automatically author scenes with diversified crowds from text input.
+To that end, we propose a novel model architecture based on dual Variational Graph Autoencoders (VGAEs) [13] that can learn how agents interact with each other and how they behave with respect to their environment.
+The trained model can be conditioned mainly on natural language, allowing users to automatically author scenes with diversified crowds from text input.
 Our approach is inspired by recent trends in generative AI.
@@
   1. 1.
-    Gen-C, a framework for text-guided authoring of high-level crowd behaviors that
+    Gen-C, a framework for text-guided authoring of high-level crowd behaviors, that
   2. 2.
-    breaks free from the dependency on real-world crowd data via the use of LLMs and simulations through
+    breaks free from the dependency on real-world crowd data via the use of LLMs and simulations, through
   3. 3.
-    a text-conditioned graph generative model that relies on a dual VGAE architecture to encode the spatial and temporal evolution of agent behaviors, interactions, and movement.
+    a text-conditioned graph generative model which relies on a dual-VGAE architecture to encode the spatial and temporal evolution of agent behaviors, interactions, and movement.
 2 Related Work
@@
 Early approaches to authoring large crowds included the use of pre-computed collision free trajectories in the form of crowd patches [17, 18] or crowd animations that could be connected together to form larger simulations  [19].
-Others, devised user interactive tools to allow for the definition of navigation fields in pre-defined environments [20].
+Others devised user interactive tools to allow for the definition of navigation fields in pre-defined environments [20].
 For authoring navigation behavior in crowd scenes, past literature heavily uses sketch-based interfaces [21, 22].
@@
 For example, Kwon et al.
-[25], propose deformation gestures as a post-processing tool for group motion editing.
+[25] propose deformation gestures as a post-processing tool for group motion editing.
 One of the most challenging aspects in authoring is controlling high-level behaviors such as describing agendas, desires and even personalities; these behaviors benefit the expressiveness and realism of simulations [26].
@@
 Data-driven approaches have also been prominent in the literature, where the crowd simulation model is implicitly defined by example data.
-Interestingly, early data-driven approaches were graph-based [33, 25], similar to early methods in the character animation literature [34].
+Interestingly, early data-driven approaches were graph-based [33, 25, 34], similar to early methods in the character animation literature [35].
 These methods however remain limited to group navigation (e.g., flocking) and fail to reflect the variations of behaviors found in human crowds.
-The need to simulate behavior diversity found in real-life crowds, led to more sophisticated and practical approaches, including approaches that learn crowd models from trajectories extracted from real-world crowd data [35, 36, 37], including works that define secondary high-level actions for increased realism [38, 39], blending different crowd styles represented by the input data [21], and more recent solutions like deep reinforcement learning [40], and even in combination with imitation learning [16, 11].
+The need to simulate the behavioral diversity found in real-life crowds has led to more sophisticated and practical approaches.
+These include methods that learn crowd models from trajectories extracted from real-world crowd data [36, 37, 34], as well as works that define secondary high-level actions to increase realism [38, 39].
+Additional approaches involve blending different crowd styles represented in the input data [21], and more recent solutions make use of deep reinforcement learning [40, 41].
+Some even combine reinforcement learning with imitation learning [16, 11].
 Despite the significant improvements achieved by these methods, the results highly depend on the variability and amount of input data.
-Contrary to all of these works, Gen-C defines heterogeneity through the use of crowd scenario graphs that are automatically generated through LLM queries and a basic simulator allowing us to move away from the dependency on collecting real crowd data.
+Contrary to all of these works, Gen-C defines heterogeneity through the use of crowd scenario graphs that are automatically generated through LLM queries allowing us to move away from the dependency on collecting real crowd data.
 2.3 Large Language Models and Virtual Agents
 Large language models have shown to be effective to domains beyond natural language processing such as task planning, embodied reasoning, autonomous agents, and floor plan design [42, 43, 44, 45].
-Here, we explore an architecture that allows a LLM to generate a sequence of crowd behaviors over time and space that can then be generalized through simulations to create synthetic crowd data.
+Here, we explore an architecture that allows an LLM to generate a sequence of crowd behaviors, over time and space, that forms synthetic crowd datasets.
 We further exploit the power of generative NLP models to enable text-guided synthesis of high-level crowd behaviors.
@@
 Along these lines, the recent work of [15] uses a text-based diffusion model to steer crowds of agents but does not extend to other types of behaviors.
-Our work seek to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
+Our work seeks to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
 2.4 Graph Generation
@@
 Sequential models like in [58] utilize conditions to initialize node and edge creation, while others incorporate conditional codes into message-passing schemes [59].
-Complementary to these works , our Gen-C framework embeds both graph structures and node features into two distinct, prior-conditioned latent distributions, enabling their generation based solely on conditional text input.
+Complementary to these works, Gen-C framework embeds both graph structures and node features into two distinct, prior-conditioned latent distributions, enabling their generation based solely on conditional text input.
 3 Gen-C Overview
@@
 We define a High-Level Scenario as the description of actions, interactions, and environment structure.
-Action is defined as a label from $\mathcal{A}ct=~{}${stand still, sit, wait, wander, queue, object interact, talk, meet, enter/exit, leave group, talk to phone, wave at, read, look at, carry}, that guides the agent’s behavior.
+Action is defined as a label from $\mathcal{A}ct=$~{wait, sit, wander, queue, enter/exit, wave at, discuss, meet, object interact, talk to phone, read, look at, carry}, that guides the agent’s behavior.
 When multiple agents perform a shared action, e.g.
@@
 1): Synthetic Data Generation, Learning, and Crowd Scenarios Generation.
-To generate the synthetic data, we first use a LLM to create diverse crowd scenarios through targeted queries that capture context, environment placement, action probabilities, and movement probabilities.
-These scenarios are extended using a preliminary simulator to introduce randomness and generalize the data, which is then transformed into graph-like structures representing action sequences, relationships, and interactions.
+To generate the synthetic data, we first use a LLM to create diverse crowd scenarios through targeted queries that capture the actions and interactions of individuals and the environment placement.
+These scenarios are transformed into graph-like structures representing action sequences, relationships, and interactions.
 Given such graphs, we train two VGAE models: one learns to reconstruct the graph structures capturing agent interactions, while the other focuses on node features representing agent actions and navigation.
-A text-conditional network acts as a prior, integrating textual information to ensure the generated graphs align with given descriptions.
+A text-conditional network acts as a prior, integrating textual information to ensure the generated graphs align with the given descriptions.
 The trained models can then synthesize novel crowd scenarios by conditionally sampling from the learned latent distributions, generating contextually aligned scenarios.
@@
 4.1 Querying Crowd Scenarios
-To generate simulations that faithfully reflect crowd dynamics, we use a series of tailored LLM queries, each designed for a specific task.
+To generate simulations that faithfully reflect crowd dynamics, we use two tailored LLM queries, each designed for a specific task.
 The process begins with a single sentence input $S_{in}$ describing a crowd scenario, which acts as the initial seed for the generation of a scenario (see top-left in Fig.
 1 for an example input).
-Specifically, we select OpenAI’s gpt-4o [59] model and run four sequential queries as presented below:
-Expanding the Scenario ($Q1$).
-The first query requests the LLM to expand $S_{in}$ into a detailed paragraph $Q1_{par}$ that describes the scenario.
-This enriched description serves as the foundation for further steps as it includes individual agent behaviors and interactions, while also provides information about the environment context.
-Defining the Environment ($Q2$).
-Using the detailed description $Q1_{par}$ and a randomly selected environment size, the second query generates a meaningful environment layout $Q2_{env}$ specifying locations and areas relevant to the scenario.
-For example, the output might define places such as a “coffee shop”, “park”, or “entrance area”, along with their position, scale and orientation.
-Assigning Location-Specific Actions ($Q3$).
-For each generated location in $Q2_{env}$, the third query identifies a list of plausible actions from $\mathcal{A}ct$ that agents might perform in that area.
-For instance, in a “coffee shop,” actions like “queue,” “meet,” or “sit” are expected.
-The LLM takes as input the $Q1_{par}$ and $Q2_{env}$ and is requested to select the top-5 actions that are most suitable for each area and assign selection probabilities to them; we ensure the total probability sums to 1.
-We note that we allow multiple possible actions per location to promote diversity in the generated simulations.
-Building Movement Plans ($Q4$).
-The fourth query focuses on defining how agents move between different locations within the environment.
-This step is similar to $Q3$, however, assigns movement probabilities from one location to every other in the environment, creating a dynamic movement plan for the agents.
-Finally, we combine the responses from all four queries and form a high-level crowd scenario $S$, generated by a high-level textual description $S_{in}$.
-4.2 Collecting Individual Behaviors
-After gathering multiple crowd scenarios from the LLM, we need a mechanism to utilize them so we can generate individual actions and interactions for each agent.
-Thus, we construct a simple simulator in Unity, where we parse each scenario $S$ and run multiple simulations.
-For each LLM scenario, we build the environment as described by $Q2_{env}$.
-Then, we randomly spawn agents in the environment, either as individuals or in small groups of two or three (all agents in a group perform the same action).
-Initially, each agent or group, queries its current location action list and randomly selects a future action based on the assigned probabilities; each action has its own duration based on the context.
-After completing the action, the agent or group of agents selects the next location based on the movement probabilities of its current location.
-This process repeats in cycles, with data collected over a period of 2 to 3 minutes.
-To introduce variability and prevent agents from consistently selecting actions and movements with the highest probabilities, we use a temperature parameter $\text{Temp}\in[.7,1]$ to amplify lower probabilities.
-Temp is randomly initialized and gradually increased to 1 by the end of the simulation.
-Finally, for each agent, we record its action sequence throughout the simulation.
-Each entry in the sequence captures the selected action and its location.
-In case the current action is shared with others (interaction), we include the ID of other agent/s in the entry.
-This detailed record $Rec_{\alpha_{i}}$ for each agent ${\alpha_{i}}$, provides a comprehensive view of the agent’s behavior and interactions within the simulated environment over its entire lifespan.
+Specifically, we select OpenAI’s gpt-4.1 [60] model and run two sequential queries as presented below:
+Generating the Environment ($Q1$).
+The first query requests the LLM to generate a meaningful environment layout, using $S_{in}$ as input, specifying locations and areas relevant to the scenario.
+For example, the output might define places such as a “coffee shop”, “park”, or “entrance area”, along with their position and scale.
+Authoring Crowd Events ($Q2$).
+The second query focuses on defining the action sequence and interactions of agents, both with the environment and other agents.
+Specifically, this request takes as input $S_{in}$ and a simplified list of the previously generated environment.
+The output is a list of multiple crowd events.
+Each event contains a sequence of action-location pairs and the id of the agent/s that participate in the current event.
+Additionally, a trigger field defines the start condition of each event, e.g., E8 starts once E2 ends.
+Finally, we combine the responses from both queries to construct a high-level crowd scenario $S$, based on a textual description $S_{in}$.
+For each agent $\alpha_{i}$} we record the sequence of events in which it appears and define a record $Rec_{\alpha_{i}}$.
+This record offers a comprehensive view of the agent’s behavior and interactions throughout its entire lifespan in the simulated environment.
 4.2 Crowd Scenario Graph Representation
@@
 In particular, we introduce the concept of crowd scenario graph, which is a a time-expanded graph that encodes both the actions and interactions of agents over a fixed temporal horizon, modeling the dynamic behaviors and relationships inherent in crowd scenarios.
-Figure 2: Structure of a crowd scenario graph. Each node contains an action label, a location category, an agent ID, and timestep as $\alpha_{i}$_$t$.
+Figure 2: Structure of a crowd scenario graph. Each node contains an action label, a location category, an agent ID, and timestep as $\alpha_{i}$_tt.
 As described in Section 4.1, during simulation, for each agent $\alpha_{i}$}, we collect $Rec_{\alpha_{i}}$.
@@
 5 Learning Crowd Scenario Graphs
-While a synthetic crowd scenario graph can be used to drive a crowd of simulated agents, ideally we want to be able to generalize without having to manually compute a new graph from scratch every time the input conditions change.
-To do so, given a dataset consisting of crowd scenario graphs, we will employ a generative model that can reason about the graphs and generate new ones.
+While a synthetic LLM-generated crowd scenario graph can be used to drive a crowd of simulated agents, ideally we want to be able to generalize without having to manually query an LLM to compute a new graph from scratch every time the input conditions change.
+To do so, given a dataset consisting of crowd scenario graphs, we will train a generative model to generated novel graphs.
 5.1 Data Preprocessing
@@
 The adjacency matrix $\mathcal{A}_{G_{i}}\in\mathbb{R}^{n\times n}$ of a graph $G_{i}$ is a symmetric matrix that encodes edge information in the graph.
-The value of the $i^{th}$ row and $j^{th}$ column is equal to 00 if there is no edge between $v_{i}$ and $v_{j}$.
-Otherwise, it set to $1$ or $-1$, if edge’s type is “sequence” or “share” respectively.
+The value of the $i^{th}$ row and $j^{th}$ column is equal to 0 if there is no edge between $v_{i}$ and $v_{j}$.
+Otherwise, it set to $1$ or 1-1, if edge’s type is “sequence” or “share” respectively.
 Node Ordering.
@@
 Node Features.
-We annotate the graphs nodes using feature vectors.
+We annotate the graph nodes using feature vectors.
 The feature vector of each node includes: the action, location, unique agent id, sequence index and interaction status.
-First, we use action and location as two one-hot vectors (sizes 15 and 8) from the available actions $\mathcal{A}ct$ and available location categories $\mathcal{L}oc$.
-For agent ID (${\alpha_{i}}$) and timestep $t$ in sequence, we set the maximum number of agents per subgraph equal to 6 and the maximum sequence length to 10; these choices reduce the complexity of generated graphs.
+First, we use action and location as two one-hot vectors (sizes 13 and 8) from the available actions $\mathcal{A}ct$ and available location categories $\mathcal{L}oc$.
+For agent ID ($\alpha_{i}$}) and timestep $t$ in sequence, we set the maximum number of agents per subgraph equal to 50 and the maximum sequence length to 45; these choices reduce the complexity of generated graphs.
 We construct another two one-hot encodings (sizes 6 and 10) for them too.
@@
 To generate the final node features, we use embedding layers to ensure that node features are dense and learnable representations suitable for downstream tasks.
-Each encoding is then passed through its respective embedding layer: actions ($\mathbb{R}^{15\times 6}$), locations ($\mathbb{R}^{8\times 4}$), agents ($\mathbb{R}^{6\times 5}$), timestep ($\mathbb{R}^{10\times 4}$), and interaction status ($\mathbb{R}^{2\times 2}$).
+Each encoding is then passed through its respective embedding layer: actions ($\mathbb{R}^{13\times 5}$), locations ($\mathbb{R}^{8\times 4}$), agents ($\mathbb{R}^{6\times 8}$), timestep ($\mathbb{R}^{4}$ using sinusoidal encoding), and interaction status ($\mathbb{R}^{2\times 2}$).
 The embeddings are concatenated to form the final node feature representation $\mathbf{X}\in\mathbb{R}^{n\times d}$ where $d$ is the feature dimension size.
 Additionally, we include the top 4 eigenvectors of the normalized graph Laplacian as node features to encode global structural properties [61].
-Thus, the final node feature dimension becomes $d=21+4=25$.
-Textual Condition.
-We convert $S_{in}$ to textual condition by first detecting and emphasizing verbs in the text to highlight action-relevant information; we utilize the spaCy library [61].
-Each $S_{in}$ is augmented by prefixing it with a list of extracted verbs, and its embedding is generated using the pretrained Sentence Transformer model all-MiniLM-L6-v2 [62].
+Thus, the final node feature dimension becomes $d=25$.
+Textual Condition. We convert $S_{in}$ into a textual condition by computing its embedding using the pretrained Sentence Transformer model all-MiniLM-L6-v2 [62].
 Figure 4: Training Model Architecture.
@@
 Graph Structure Decoder.
-Reconstructs the padded graph adjacency matrix $\mathcal{A}_{G^{\prime}}$ from $Z_{S}$; we set the maximum number of nodes equal to 40.
+Reconstructs the padded graph adjacency matrix $\mathcal{A}_{G^{\prime}}$ from $Z_{S}$; we set the maximum number of nodes equal to 50.
 It uses a MLP to transform the concatenated latent and conditional representations into a flattened adjacency matrix.
@@
 Each GINEConv layer applies a message-passing operation, followed by non-linear transformations using LeakyReLU activations and LayerNorm for stability.
-The final layer projects the hidden representations into the node feature space, with size 23.
+The final layer projects the hidden representations into the node feature space, with size 21.
 Condition Net.
-Both VGAE models use the same condition network structure that encodes the embeddings of text ($S_{in}$) and number of agents to produce the condition vector $C$.
+Both VGAE models use the same condition network structure that encodes the embeddings of text ($S_{in}$), number of agents, and number of nodes to produce the condition vector $C$.
 We utilize $\mathbf{C}$ to train two Prior Networks that parameterize the prior distributions $p(Z|C)$ of the latent variables of VGAE-S and VGAE-F, computing their respective means $\mu_{\text{prior}}$ and standard deviations $\sigma_{\text{prior}}$.
@@
 $\mathcal{L}_{\text{KL}}=\mathbb{E}_{z\sim q(Z)}\left[\log q(Z\mid X,C)-\log p(Z\mid C)\right].$ (3)
-[…]
+We use a scale factor of $\beta\in[0,4]$ to balance reconstruction losses $\mathcal{L}_{S}$ and $\mathcal{L}_{F}$ against $\mathcal{L}_{KL}$ and employ a cyclical annealing schedule [64], periodically ramping $\beta$ from 0 its maximum.
+The final training loss for the two VGAE models are: $\mathcal{L}_{VGAE-G}=\mathcal{L}_{S}+\beta\mathcal{L}_{KL_{S}}$, and $\mathcal{L}_{VGAE-F}=\mathcal{L}_{F}+\beta\mathcal{L}_{KL_{F}}$.
 6 Text-Conditioned Scenario Generation
@@
 We parse the generated graphs to collect a sequence of actions-locations and interactions per agent.
-Then, to generate a simulation, a simulation environment can be either (a) manually provided by the user or (b) generated with a simple LLM query similar to $Q2$, as described in Section 4.1.
+Then, to generate a simulation, a simulation environment can be either (a) manually provided by the user or (b) generated with a simple LLM query similar to $Q1$, as described in Section 4.1.
 Given the environment, we run simulations of the generated scenarios in Unity [5].
@@
 For each dataset, we begin by defining 100 unique input sentences describing crowd scenarios; an example input is “Outside the station, passengers sit on benches to read a newspaper and talk on the phone.” Using the process detailed in Section 4.1, we generate 5 variations for each sentence, resulting in 500 crowd scenarios per dataset.
-Graph datasets are then constructed as outlined in Section 4.2 and Section 4.3, with the University Campus dataset containing $135k$ subgraphs and the Train Station dataset $110k$.
+Graph datasets are then constructed as outlined in Section 4.2, with the University Campus dataset containing $12.5k$ subgraphs and the Train Station dataset $13.8k$.
 The training data and models will be released upon acceptance.
@@
 The graph structure decoder uses 3 linear layers with a hidden dimension of 128, applying a tanh activation after the final layer.
-The feature decoder employs 3 GINE layers with a hidden dimension of 128, followed by a layer mapping to a 23-dimensional feature space.
+The feature decoder employs 3 GINE layers with a hidden dimension of 128, followed by a layer mapping to a 21-dimensional feature space.
 The condition network projects text embeddings from 384 to 128 dimensions and agent embeddings from 32 to 128, combining them into a conditional vector of size 128.
@@
 Training each model required approximately 4 hours on a single NVIDIA RTX 4070 Ti GPU.
-Quantitative Results.
+Reconstruction Performance.
 We test the reconstruction quality of our model relying solely on the conditional prior $p(Z|C)$.
@@
 For instance, in the Train Station dataset, the "carry" label is more frequent, while "wander" is less common compared to the University Campus dataset, reflecting travelers often carrying bags and students wandering in outdoor areas.
-We also evaluate the diversity of feature generation in our model.
-Using the same textual description $S_{in}$, we sample 1k variables from the latent space of VGAE-F and collect each label’s frequency.
-The results for four user-input scenarios are presented in Fig.
-6, demonstrating that the model produces diverse outputs rather than being limited to a fixed set.
-Moreover, the generated results remain consistent with the input condition $S_{in}$.
+Real Data Consistency.
+Gen-C uses the knowledge that LLMs acquired from text data to synthesize crowd scenario graphs, reducing reliance on real-world video data.
+To evaluate this approach, we annotated high-level actions in the Students03 video from the UCY dataset [37], which features 434 students over 3 minutes and 36 seconds.
+We segmented the video into 10-second windows and recorded the following action frequencies: discuss = 287, meet = 61, talk to phone = 14, carry = 15, wait = 13, sit = 52, wander = 3, and wave at = 5.
+As not all available actions are present in the real video, actions with zero frequency (e.g., read, look at, queue, enter/exit) are omitted.
+As shown in the top right of Fig.
+6, our model’s action distribution for the University Campus closely aligns with this real data, showing similarly high values for discuss and meet, and low for talk to phone and wait.
+The model overproduces sit and wander, likely due to the real scene capturing a small-medium size area lacking seating areas.
+Action Sequence Analysis.
+To analyze the sequence of actions in our data, we perform the following procedure.
+For each scenario, we gather all graphs from the dataset and count the occurrences of each subsequent action for every current action.
+The results are shown in Fig.
+5.
+We observe that action transitions are generally intuitive—for example, "discuss" frequently follows "meet", and "wander" often follows "enter/exit", which aligns with behavior when entering a new area.
+Additionally, certain actions are more prominent based on the scenario: for instance, "carry" shows higher frequencies in the Train Station, as people often carry luggage in this setting.
+Figure 5: Action Sequence Tables. We present the frequency of each future action for each action label.
+Latent Space Analysis.
+To quantify how closely our generated graphs inhabit the same learned latent manifold as the training data, we sampled a batch of graphs from the learned prior and encoded both trained and predicted graphs to obtain two sets of embeddings, $z_{\mathrm{train}}$ and $z_{\mathrm{pred}}$.
+We visualized these embeddings in Figure 8, where we observe well-mixed clusters with no clear signs of mode collapse.
+To measure the distributional gap between train and generated embeddings, we computed the Fréchet Distance (FD) [66] between their Gaussian distributions:
+$\mathrm{FD}=\|\mu_{\mathrm{train}}-\mu_{\mathrm{pred}}\|_{2}^{2}+\mathrm{Tr}\left(\Sigma_{\mathrm{train}}+\Sigma_{\mathrm{pred}}-2\sqrt{\Sigma_{\mathrm{train}}\Sigma_{\mathrm{pred}}}\right).$
+A lower FD indicates a closer alignment between distributions.
+For feature embeddings, we observe very low FD values (Train Platform: 3.6; University Campus: 6.0), suggesting near-perfect feature recovery.
+In contrast, the FD for graph structure embeddings is higher (Train Platform: 33.2; University Campus: 19.4), yet still reflects reasonable diversity without significant drift from the real-data manifold.
 Ablation Study.
@@
 w/o Cano. Single Gen-C w/o Cano. Single Gen-C
-Degree .2578 .0053 .0045 .4710 .0045 .0055
-Clustering Coeff. .5143 .0203 .0046 .4511 .0096 .0074
-Diameter .1996 .1425 .1200 .2642 .0369 .0820
-Avg. Path Length .1877 .2244 .0360 .6330 .1005 .0500
-Action .4149 .3009 .0640 .2699 .3078 .1100
-Location .0925 .0316 .0027 .0270 .0480 .0088
+Degree 1.087 .620 .590 1.120 .107 .110
+Clustering Coeff. 2.289 .015 .004 2.170 .0083 .003
+Diameter 1.326 .723 .370 1.401 .494 .330
+Avg. Path Length 1.602 .310 .150 1.632 .169 .200
+Action 0.210 .141 .130 .118 .213 .066
+Location 0.059 .032 .019 .021 .0512 .015
 8 Discussion and Future Work
@@
 Our approach leverages an LLM and two learned latent spaces to enable the generation of crowd scenario graphs that drive agent behaviors.
-Overall, our system generate plausible simulations, allowing the automatic creation of virtual environments with agents exhibiting diverse, high-level actions beyond typical low-level crowd behaviors like goal-seeking and collision avoidance.
+Overall, our system generates plausible simulations, allowing the automatic creation of virtual environments with agents exhibiting diverse, high-level actions beyond typical low-level crowd behaviors like goal-seeking and collision avoidance.
 However, several limitations exist.
-The simulator used to expand LLM-generated scenarios relies on probabilistic action sequences, thus the sequence of actions may not be optimal.
+Our current implementation does not support modeling long-term agent intentions.
 Agents cannot switch actions mid-execution, and action durations are manually defined.
 Additionally, behaviors are constrained to a predefined list of high-level actions, requiring retraining to accommodate new actions.
-For future work, we aim to integrate our framework with a crowd simulation framework, allowing high-level behaviors to control low-level navigation policies, creating more dynamic and realistic simulations.
+For future work, we aim to equip agents with a memory module, ensuring a more natural selection and progression of actions.
+We would also like to combine Gen-C with a crowd simulation framework, allowing high-level behaviors to control low-level navigation policies, creating more dynamic and realistic simulations.
 Additionally, we plan to explore combining latent spaces from different datasets, such as "indoor" and "outdoor" scenarios, using techniques such as distillation to enable lifelong learning and dynamic simulation adaptability.